OpenStudy (anonymous):

Write an equation for the lines with the given information. (In the comments, four separate questions) PLEASE HELP?!

OpenStudy (anonymous):

a. passes through (-1,4) and (-5,2) b. slope 2, y-intercept -4 c. has an x-intercept of 6 and a y-intercept of 3 d. passes though (1,2) with slope -5/3

OpenStudy (lstrasmann):

b. \[ y = 2x - 4\] d. \[ y - 2 =-\frac{ 5 }{ 3 }(x - 1)\]

OpenStudy (lstrasmann):

Wait, I just remembered how to do A. \[\frac{ 2-4 }{ -5+1 } = \frac{ -2 }{ -4 } = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Thank you so much! ^-^